OpenStudy (asib1214):

A glass jar with a volume of 215cm^3 is designed to withstand a pressure of 253 KPa. The jar is filled with gas at a pressure of 152 KPa at ambient temperature. Can the glassware withstand the gas being heated to 2X10^2 Degree Celsius?

OpenStudy (asib1214):

Pv=nRT the variables i think i'll be working with are P/T=nR/V P1/T1=P2/V2, where V1 is 152/25=253/T

OpenStudy (somy):

i suppose its better to use P1/T1 = P2/T2

OpenStudy (asib1214):

thats what i'm using but for some reason i'm not getting the right answer

OpenStudy (asib1214):

sorry if i mistyped the furmula but that;s exactly what i'm using to find the temperature

OpenStudy (somy):

alright did u do it like this? 253 000/ X = 152 000/ ((2X10^2) + 273)

OpenStudy (asib1214):

i get, T2= 253X25/152=41.61C

OpenStudy (somy):

look what i did

OpenStudy (asib1214):

sorry for the late reply, i think the OS went down right after replied to my comment. Anyways why did you multiply the pressure by 1000????

OpenStudy (somy):

coz its in Kpa i need in Pa

OpenStudy (somy):

SI system units

OpenStudy (asib1214):

if i calculate the way you stated above, i get 787.30 which is not right

OpenStudy (somy):

787.30- 273 =

OpenStudy (asib1214):

514.3, which is still not right :(

OpenStudy (somy):

what is the right answer?

OpenStudy (asib1214):

the answer is 810 or 8.1X10^2

OpenStudy (asib1214):

8.1X10^2

OpenStudy (somy):

thats so weird

OpenStudy (asib1214):

yep!!!! wait let me try myself

OpenStudy (asib1214):

i extremely apologize the right answer is 241KPa and not 810...... i've double checked this time....

OpenStudy (asib1214):

@Somy

OpenStudy (somy):

oh

OpenStudy (asib1214):

this is how i tried to do the questioin, 152X(101.3+273)/253 and i get 224.8KPa

OpenStudy (somy):

101.3 ? where did u get this from?

OpenStudy (asib1214):

it says that the experiment was done in 1 atm pressure which is 101.3

OpenStudy (asib1214):

or ambient temperature

OpenStudy (asib1214):

25C

OpenStudy (somy):

oh i see btw if u r saying that 101.3 is pressure then u r doing the calculation wrong

OpenStudy (asib1214):

yah i jus realized LOL

OpenStudy (asib1214):

i have no iead what i'm suppose to do....

OpenStudy (somy):

lol 273 + Celsius degree = Kelvin its unit conversion

OpenStudy (somy):

u are adding 273 to pressure which is obviously wrong

OpenStudy (asib1214):

so what do i do then?????

OpenStudy (somy):

first tell me what formula u used just now? im confused

OpenStudy (asib1214):

P1/T1 = P2/T2, where P1 is the initial pressure 152KPa/initial 25C= 253KPa/T2

OpenStudy (somy):

but then you will find temperature tho

OpenStudy (asib1214):

, the question asks whether or not the temperature i get is higher or lower than 200C which is what the glassware is designed for. so i get get i number less than 200, that means that the glassware can withstand the heat if not the temperature will be higher than 200C

OpenStudy (somy):

im getting 223 Celsius

OpenStudy (asib1214):

i'm getting 224

OpenStudy (somy):

you are confusing me u said i extremely apologize the right answer is 241KPa and not 810...... i've double checked this time.... kpa is PRESSURE

OpenStudy (somy):

so we need pressure

OpenStudy (asib1214):

yah i jus realized that too.... so what am i suppose to do :'(

OpenStudy (asib1214):

set the equation so that i'm solvign for pressure????

OpenStudy (somy):

ok let me think properly coz i got too confused ~_~ with what u r saying and what the question is saying

OpenStudy (somy):

got it

OpenStudy (somy):

right

OpenStudy (asib1214):

limit of the pressure of the jar = 253 KPa @ ambiant/25C intial pressure of the jar= 152 KPa what would be the pressure of the jar if the temperature is 200C

OpenStudy (somy):

look we will use P1/T1 = P2/T2 P1= 152 Kpa T1= 25 Celsius +273= 298 K P2= X T2= 200 + 273= 473 K

OpenStudy (somy):

so 152/298 = X/ 473 X= 473 *152/ 298 = 241kPa

OpenStudy (somy):

right now i get the question ~_~ u mixed up the whole conversion stuff so i got confused

OpenStudy (somy):

look when we use Gas law formulas we have to use temperature in Kelvin always remember that you used Celsius thats why you were getting stuff wrong + well you got the question wrong a bit too but well you were on the right way so basically saying the question asked if it will with stand - if the temperature is increased now that means we need to find pressure to which this temperature will lead us to, and then compare

OpenStudy (asib1214):

i really appreciate the effort you put into solving the question. I wish i could give you a noble prize but for now thank you so much!!!!!!!!

OpenStudy (somy):

lol you are welcome :)))))